OpenStudy (mikey):

medal and fan for answer Simplify 9 over quantity of 3 minus 2 i.

OpenStudy (mikey):

@3mar @AloneS

OpenStudy (3mar):

Can you write it? I know how to write it, just want you to train your fingers! ;)

OpenStudy (3mar):

\[\frac{ 9 }{ 3-2i }\] we would multiply the numerator and denominator by (3+2i)

OpenStudy (3mar):

let's try \[\frac{ 9 }{ 3-2i }*\frac{ 3+2i }{ 3+2i }=\frac{ 9(3+2i) }{ 9+4 }=\frac{ 9(3+2i) }{ 13 }\]

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (3mar):

Hope that helps